OpenStudy (anonymous):

Let u = <5, 6>, v = <-2, -6>. Find -2u + 5v.

OpenStudy (anonymous):

The answers are: <-20, -42> <-20, 0> <0, 18> <-6, 0>

OpenStudy (anonymous):

I have completely forgotten how to do this. Any help? @phi

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (phi):

-2u means -2 times each component of u ditto 5v then add "corresponding" components

OpenStudy (anonymous):

So for u: 5 * -2 * -2? For v: 5 * 6 * -6? I don't see any answers that correspond with that? @phi

OpenStudy (phi):

-2u means -2*<5, 6> = <-2*5, -2*6>= <-10,-12>

OpenStudy (anonymous):

Nevermind I see what I did wrong.

OpenStudy (anonymous):

It's 5 * -2, -2 * 6 = -10, -12 and 5 *6, 5 * -6 = 30, -30 so it's -20, -42?